# Assessing Spatial Disparities: A Bayesian Linear Regression Approach

### v4

@@ 1. Introduction
We develop statistical methods for detecting spatial disparities on maps. Disparities in health outcomes are a rapidly evolving field within public health (see, e.g., rao2023) and spatial variation in health outcomes affects how disparities manifest on disease maps. Spatial data analysis in epidemiological investigations is extensively documented (see, e.g., waller2004applied; waller2010handbook; lawson2013statistical; lawson2016handbook, and references therein). Inference on disparities proceeds from “difference boundaries” that represent significant differences between neighbors and delineate regions with significantly different disease mortality and incidence rates, helping to improve the geographic allocation of health care resources. In geographic analysis, the problem is often called “wombling” (see jacquez2003local; jacquez2003geographic; lu2005bayesian; li2011mining; ma2007bayesian; fitzpatrick2010, for some algorithmic approaches with applications). Probabilistic model-based approaches include developments in lu2007bayesian; ma2010hierarchical; li2015bayesian; banerjee2012bayesian; hanson2015spatial; leeBoundaryDetectionDisease2012; corpas-burgos2020serra; gaoSpatialDifferenceBoundary2023 and aiello2023detecting.
@@
The aforementioned articles present an impressive range of methods, but are broadly classified into two groups. The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements. The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities. Both groups introduce spatial dependence and aim to balance the underlying spatial similarities with the ability to detect differences among neighbors. Our current contribution offers a computationally scalable and analytically tractable approach to this problem. Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (mullerFDRBayesianMultiple2006). The benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects of adjacent regions in determining spatial disparity, and providing improved sensitivity and specificity for the detection of boundary differences (as seen in the analysis by gaoSpatialDifferenceBoundary2023). Our parametric approach is suitable for applications where one disease map is of interest and lack of replication hinders non-parametric methods (such as in slackGeographicConcentrationUs2014; jagai2017county; rodriguez2018county).
@@
We exploit closed-form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold. We extend lu2005bayesian and fitzpatrick2010, who use only posterior means or spatial residuals to detect differences, to fully model-based inference. We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model. Section LABEL:sec:simulation_experiments evaluates our methods alongside an existing Bayesian nonparametric approach using simulated data over a map of Californian counties followed by an analysis of lung cancer mortality data in Section LABEL:sec:application. Section LABEL:sec:_discussion concludes with a brief discussion.
@@ 2.1. Random Effects under an Augmented Linear System
We follow rieblerIntuitiveBayesianSpatial2016 who adapted the “BYM” model (besagBayesianImageRestoration1991) to ensure identifiable parameters and provide a better interpretation. We write this “BYM2” model as
$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim N_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim N_{n}(0,I_{n})$ (1)
where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix such that $\pi(\phi)$ is proper. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is equal to one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (rieblerIntuitiveBayesianSpatial2016). For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, one of the most popular spatial models used in areal data analysis. Here, $\pi(\gamma_{i}\,|\,\sigma^{2},\gamma_{j},j\neq i)=N(\alpha\sum_{j}w_{ij}\gamma_{j}/w_{i+},c\sigma^{2}/w_{i+})$, where $w_{ij}$ is the binary adjacency indicator for regions $i$ and $j$, $\alpha$ is a spatial smoothing parameter, $c$ is a fixed scaling constant and $w_{i+}=\sum_{j}w_{ij}$ is the number of neighbors of region $i$. The conditional distribution of each spatial residual is based on a linear combination of its neighbors. By Brook’s lemma (besagBayesianImageRestoration1991), this equates to setting $V^{-1}_{\phi}=c(D_{W}-\alpha W)$ in (1), where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}=w_{i+}$ (banerjeeHierarchicalModelingAnalysis2015, Chap. 4). The density $\pi(\phi)$ is proper when $\alpha\in(1/\lambda_{min},1)$, where $\lambda_{min}$ is the smallest eigenvalue of $D_{W}^{-1/2}WD_{W}^{-1/2}$. Since $\rho$ already controls the amount of spatial smoothing in (1), we set $\alpha=0.99$ in the following sections to improve the identifiability of spatial effects $\phi$, which we consider as potential indicators of latent factors driving local health inequity.
If $\pi(\gamma,\beta,\sigma^{2},\rho)=N(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times N(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
@@
where $V_{y_{\star}}=\begin{bmatrix}(1-\rho)I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho V_{\phi}\end{bmatrix}$ and a prior $\pi(\gamma_{\star},\sigma^{2},\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ (since the prior for $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ has been absorbed into the linear system). The resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y,\rho)$, is
@@
where $M_{\star}=(X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star})^{-1}$, $m_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}=\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}X+m_{0}^{\mathrm{\scriptscriptstyle T}},\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}\right)^{\mathrm{\scriptscriptstyle T}}$, $a_{n}=a_{\sigma^{2}}+\frac{n}{2}$, and $b_{n}=b_{\sigma^{2}}+\frac{1}{2}\left(\frac{1}{1-\rho}y^{\mathrm{\scriptscriptstyle T}}y+m_{0}^{\mathrm{\scriptscriptstyle T}}M_{0}m_{0}-m_{\star}^{\mathrm{\scriptscriptstyle T}}M_{\star}m_{\star}\right)$. The posterior mean $M_{\star}m_{\star}=\widehat{\gamma}_{\star}$ is the general least squares regression estimate to be solved $X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}X_{\star}\widehat{\gamma}_{\star}=X_{\star}^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}y_{\star}$ and $b_{n}=b_{\sigma^{2}}+\frac{ns^{2}}{2}$, where $s^{2}=\frac{1}{n}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})^{\mathrm{\scriptscriptstyle T}}V_{y_{\star}}^{-1}(y_{\star}-X_{\star}\widehat{\gamma}_{\star})$ is the mean residual sums of squares from (2).
If $\rho\in(0,1)$ is fixed, then we can draw exact posterior samples from (3). Noting that $\gamma=\sigma\sqrt{\rho}\phi$ and carrying out some familiar linear algebra calculations, we obtain
@@
These expressions are used in Section 3.2 to formulate the detection of differences between boundaries.
@@ 2.2. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$. By spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are nonnegative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}$, and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y$. Calculations reveal $\lim_{\rho\rightarrow 0^{+}}\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)=0_{n\times n}$ and $\lim_{\rho\rightarrow 0^{+}}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]=0_{n}$. Therefore, if the prior on $\beta$ is flat, then $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)$ collapses to the degenerate mass at $\gamma=0_{n}$.
We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with the entries $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Further simplifications reveal that $\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]\rightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right)$ and $\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\rightarrow UD^{\star}U^{-1}e$ as $\rho\rightarrow 1^{-}$, where $e=(I-H)y$ are the residuals of ordinary least squares. Furthermore, $\operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}$, $\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\rightarrow\sigma^{2}V$ and $\operatorname{Cov}(\beta,\gamma\,|\,y,\sigma^{2},\rho)\rightarrow-\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}V$ as $\rho\rightarrow 1^{-}$, where $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$. Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$. These limiting cases have implications for subsequent boundary detection (see Section LABEL:subsec:epsilon_intro).

### v5

@@ 1. Introduction
We develop statistical methods for detecting local spatial disparities on maps. Disparities in health outcomes are a rapidly evolving field within public health (Rao, 2023) and spatial variation in health outcomes affects how disparities manifest on disease maps. Spatial data analysis in epidemiological investigations is extensively documented (Waller and Gotway, 2004; Waller and Carlin, 2010; Lawson, 2013; Lawson et al., 2016). Inference on disparities proceeds from “difference boundaries” that represent significant differences between neighbors and delineate regions with significantly different disease mortality and incidence rates, helping to improve the geographic allocation of health care resources and guide future investigation into local determinants. Boundary detection enables precise identification of health event clusters and sudden changes in latent or observed environmental exposures, which is crucial for defining sampling protocols across a diverse range of health environments and planning targeted interventions (Copeland, 2010; Jacquez, 2010). For example, Barboza-Salerno et al. (2025) apply boundary detection to identify local socioeconomic inequalities and estimate their association with neighborhood child mortality. In geographic analysis, the problem is often called “wombling” (see Jacquez and Greiling, 2003aa, 2003bb; Lu and Carlin (2005); Li et al. (2011); Ma and Carlin (2007); Fitzpatrick et al. (2010) for some algorithmic approaches with applications). Probabilistic model-based approaches include developments in Lu et al. (2007); Ma et al. (2010); Li et al. (2015, 2012); Hanson et al. (2015); Lee and Mitchell (2012); Corpas-Burgos and Martinez-Beneito (2020); Gao et al. (2023) and Aiello and Banerjee (2025).
@@
The aforementioned articles present an impressive range of methods, but are broadly classified into two groups. The first embraces hierarchical models and Bayesian inference for the spatial adjacency matrix by modeling its elements. The second endows the spatial random effects with a probability law that accommodates nonzero probabilities on neighboring spatial random effects being equal (often using an areal Dirichlet process or some adaptation thereof), which are computed as posterior probabilities. Both groups introduce spatial dependence and aim to balance the underlying spatial similarities with the ability to detect differences among neighbors. Our current contribution offers a computationally scalable and analytically tractable approach to this problem. Instead of modeling spatial effects as realizations of a Dirichlet process, which struggles to learn from the data and impedes convergence of iterative estimation algorithms, we devise fully model-based Bayesian inference within an augmented spatial linear regression framework. More specifically, we treat spatial disparities as a multiple comparison problem within a Bayesian inferential paradigm by testing for differences in spatial random effects between neighboring regions. Subsequently, we use posterior probabilities of such differences to detect spatial disparities while controlling a Bayesian false discovery rate or FDR (Müller et al., 2007). Controlling the false discovery rate is preferred to controlling the family wise error rate or false positive rate because it offers increased power and flexibility (Catelan and Biggeri, 2010; Glickman et al., 2014) and has been previously applied in analyses of health disparities (Tian et al., 2010; Slutske et al., 2023). Further developments of Bayesian FDR for multiple testing in spatial settings include Ventrucci et al. (2011); Sun et al. (2015); Tansey et al. (2018).
Benefits of our approach include clearer interpretation of spatial effects, computational simplicity, accounting for small differences in the values of neighboring spatial effects in determining spatial disparity, and providing improved sensitivity and specificity for detecting boundary differences (as seen in Gao et al., 2023). Our parametric approach is suitable for applications where one disease map is of interest and lack of replication hinders non-parametric methods (such as in Slack et al., 2014; Jagai et al., 2017; Rodriguez et al., 2018).
@@
We exploit closed-form distribution theory to account for the differences between the values of the spatial effects and evaluate posterior probabilities of such differences exceeding a certain threshold. We extend Lu and Carlin (2005) and Fitzpatrick et al. (2010), who use only posterior means or spatial residuals to detect differences, to fully model-based inference. We provide an outline for the rest of the article. Section 2 discusses the problem of detecting differences between fixed effects in a linear model before we extend to spatial random effects. Section 3 develops the notion of a difference boundary, demonstrates a meaningful connection to traditional hypothesis testing in the simple linear regression model, and considers extensions to a generalized linear model. Section 4 evaluates our methods alongside an existing Bayesian nonparametric approach using simulated data over a map of Californian counties followed by an analysis of lung cancer mortality data in Section 5. Section 6 concludes with a brief discussion.
@@ 2.1. Random Effects under an Augmented Linear System
We follow the adaptation by Riebler et al. (2016) of the “BYM” model (Besag et al., 1991) to ensure identifiable parameters and provide better interpretation. The “BYM2” model is
$y=X\beta+\gamma+\eta,\;\;\gamma=\sigma\sqrt{\rho}\phi,\;\;\phi\sim\mbox{N}_{n}(0,V_{\phi}),\;\;\eta=\sigma\sqrt{1-\rho}v,\;\;v\sim\mbox{N}_{n}(0,I_{n})$ (1)
where $\phi$ is an $n\times 1$ vector of random effects, $V_{\phi}$ is a known positive definite matrix such that $\pi(\phi)$ is proper. The parameter $\sigma^{2}$ captures the overall error variance if $c$ is set so that the geometric mean of the marginal prior variances is equal to one, and $\rho$ represents the proportion of total variance attributed to the spatial structure (Riebler et al., 2016). For subsequent analysis, we specify a conditionally autoregressive (CAR) structure, one of the most popular spatial models used in areal data analysis. Here, $\pi(\gamma_{i}\,|\,\sigma^{2},\gamma_{j},j\neq i)=\mbox{N}(\alpha\sum_{j}w_{ij}\gamma_{j}/w_{i+},c\sigma^{2}/w_{i+})$, where $w_{ij}$ is the binary adjacency indicator for regions $i$ and $j$, $\alpha$ is a spatial smoothing parameter, $c$ is a fixed scaling constant and $w_{i+}=\sum_{j}w_{ij}$ is the number of neighbors of region $i$. The conditional distribution of each spatial residual is based on a linear combination of its neighbors. By Brook’s lemma (Besag et al., 1991), this equates to setting $V^{-1}_{\phi}=c(D_{W}-\alpha W)$ in (1), where $W$ is the adjacency matrix, $D_{W}$ is diagonal with the diagonal elements $D_{W,ii}=w_{i+}$ (Banerjee et al., 2015, Chap. 4). The density $\pi(\phi)$ is proper when $\alpha\in(1/\lambda_{min},1)$, where $\lambda_{min}$ is the smallest eigenvalue of $D_{W}^{-1/2}WD_{W}^{-1/2}$. Since $\rho$ already controls the amount of spatial smoothing in (1), we set $\alpha=0.99$ in the following sections to improve the identifiability of spatial effects $\phi$, which we consider as potential indicators of latent factors driving local health inequity.
If $\pi(\gamma,\beta,\sigma^{2},\rho)=\mbox{N}(\gamma\,|\,0,\sigma^{2}\rho V_{\phi})\times\mbox{N}(\beta\,|\,M_{0}m_{0},\sigma^{2}M_{0})\times\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ is the prior for (1), then the joint distribution arises from the augmented linear model,
@@
where $V_{y_{\star}}=\begin{bmatrix}(1-\rho)I_{n}&0&0\\ 0&M_{0}&0\\ 0&0&\rho V_{\phi}\end{bmatrix}$ and a prior $\pi(\gamma_{\star},\sigma^{2},\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}},b_{\sigma^{2}})\times\pi(\rho)$ (since the prior for $\gamma_{\star}=(\beta^{\mathrm{\scriptscriptstyle T}},\gamma^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ has been absorbed into the linear system). By conjugacy of the prior $\pi(\gamma_{\star},\sigma^{2}\,|\,\rho)$ with the Gaussian likelihood in (2), the resulting conditional posterior joint distribution, $\pi(\gamma_{\star},\sigma^{2}\,|\,y_{\star},\rho)\propto\pi(y_{\star}\,|\,\gamma_{\star},\sigma^{2},\rho)\times\pi(\gamma_{\star},\sigma^{2}\,|\,\rho)$, is
@@
[…]
Fixing $\rho\in(0,1)$ we can draw exact posterior samples from (3). Since $\gamma=\sigma\sqrt{\rho}\phi$, we obtain
@@
Section 3.2 uses these expressions to formulate the detection of differences between boundaries.
@@ 2.2. Limiting cases
The limits of the conditional posterior distribution of $(\beta,\gamma)$ as $\rho\rightarrow 0^{+}$ or as $\rho\rightarrow 1^{-}$ are analytically derived when the prior is non-informative on $\beta$. By spectral decomposition, $V^{1/2}_{\phi}(I-H)V^{1/2}_{\phi}=CDC^{\mathrm{\scriptscriptstyle T}}$ where $C$ is an orthogonal matrix and $D=\text{diag}(d_{1},\ldots,d_{n})$ with entries that are nonnegative eigenvalues of $V_{\phi}^{1/2}\left(I_{n}-H\right)V_{\phi}^{1/2}$. Let $U=V^{1/2}_{\phi}C$, then $U$ is invertible, $U^{\mathrm{\scriptscriptstyle T}}V^{-1}_{\phi}U=I_{n}$, and $U^{\mathrm{\scriptscriptstyle T}}(I_{n}-H)U=D$. The block entries of $M_{\star}$ and $\widehat{\gamma}_{\star}$ in (3) reveal that as $\rho\rightarrow 0^{+}$,
$\begin{split}\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]&=U\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}DU^{-1}y\longrightarrow 0_{n},\\ \operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)&=\sigma^{2}U\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}U^{\mathrm{\scriptscriptstyle T}}\longrightarrow 0_{n\times n}.\end{split}$ (6)
Therefore, with a flat prior on $\beta$, $\lim_{\rho\rightarrow 0^{+}}\pi(\beta,\gamma,\sigma^{2}\,|\,y,\rho)=\delta_{0_{n}}(\gamma)$(degenerate mass at $\gamma=0_{n}$).
We also obtain $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1}{\rho}I_{n}+\frac{1}{1-\rho}D\right)^{-1}=I_{n}-D^{\star}$ and $\lim_{\rho\rightarrow 1^{-}}\left(\frac{1-\rho}{\rho}I_{n}+D\right)^{-1}D=D^{\star}$ where $D^{\star}\in\mathbb{R}^{n\times n}$ is diagonal with $D_{ii}^{\star}=I(d_{i}>0)$ for $i=1,\ldots,n$. Simplifications reveal
$\begin{split}\mathbb{E}[\beta\,|\,y,\sigma^{2},\rho]&\longrightarrow(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}\left(y-UD^{\star}U^{-1}e\right),\quad\mathbb{E}[\gamma\,|\,y,\sigma^{2},\rho]\longrightarrow UD^{\star}U^{-1}e,\\ \operatorname{Var}(\beta\,|\,y,\sigma^{2},\rho)&\longrightarrow\sigma^{2}(X^{\mathrm{\scriptscriptstyle T}}X)^{-1}X^{\mathrm{\scriptscriptstyle T}}VX(X^{\mathrm{\scriptscriptstyle T}}X)^{-1},\quad\operatorname{Var}(\gamma\,|\,y,\sigma^{2},\rho)\longrightarrow\sigma^{2}V,\end{split}$ (7)
as $\rho\rightarrow 1^{-}$, where e=(IH)ye=(I-H)y are the residuals from ordinary least squares and $V=V_{\phi}-UD^{\star}U^{\mathrm{\scriptscriptstyle T}}$. Finally, $\lim_{\rho\rightarrow 1^{-}}b_{n}=b_{\sigma^{2}}+\frac{1}{2}v^{\mathrm{\scriptscriptstyle T}}D^{\star}v$, where $v=U^{-1}e$. Therefore, $\lim_{\rho\rightarrow 1^{-}}\pi(\sigma^{2}\,|\,y,\rho)=\text{IG}(\sigma^{2}\,|\,a_{\sigma^{2}}+n/2,b_{\sigma^{2}}+v^{\mathrm{\scriptscriptstyle T}}D^{\star}v/2)$. Section 3.5 discusses their implications for boundary detection.